At an altitude of h feet above the surface of the earth, the approximate distance in miles that a person can see is given by the expression d2 = 1.5h. How far can a person see if he or she is 1270 feet above the earth's surface? (Round your answer to the nearest tenth of a mile.)
Answer by checkley77(12844)

d2 = 1.5h
D^2=1.5*1,270
D^2=1,905
D=sqrt1,905
D=43.6 MI.